How would I do these? (Sorry, long question.)
Simplify and rationalize the denominator where enccessary (assume x > 0).
a) sqrt. 16,000

Answer by richwmiller(17219)

Long involved questions are allowed but these are just a bunch of similar questions strung together.
No similar problems.
sqrt(16000)
sqrt(16*100*10)=4*10*sqrt(10)=40 sqrt(10)